PLEASE SHOW ME HOW THESE PROBLEMS ARE SOLVED

STEP BY STEP WITH CLEAR EXPLANATION WOULD BE HIGHLY APPRECIATED!1. For each of the following production functions, decide if (and EXPLAIN why) it describes a constant, increasing or decreas3. Suppose that the demand curve for wheat is D(p) = 120 - 10p and the supply curve is S(p) = 10p. (a) Find the (free market)
Answer #1

1.
Let K = tK and L = tL where t is a constant and t > 1.
a. f(K,L) = K0.5L0.8
f(tK,tL) = (tK)0.5(tL)0.8 = t0.5K0.5t0.8L0.8 = t0.5+0.8K0.5L0.8 = t1.3f(K,L)
So, power of t is 1.3 which is greater than 1. Thus, there are increasing returns to scale.

b. f(K,L) = 5(K1/2 + L1/2)3
f(tK,tL) = 5[(tK)1/2 + (tL)1/2]3 = 5(t1/2K1/2 + t1/2L1/2)3 = 5[t1/2(K1/2 + L1/2)]3 = t3/25(K1/2 + L1/2)3 = t1.5f(K,L)
So, power of t is 1.5 which is greater than 1. Thus, there are increasing returns to scale.

(Note: Post one question at a time.)